# SAMix: Calibrated and Accurate Continual Learning via Sphere-Adaptive Mixup and Neural Collapse

### v1

@@ 5 Conclusion
However, a limitation exists: while SAMix improves performance and calibration, it is tailored to fixed prototype-based CL. Our experiments with dynamic prototype methods and traditional one-stage CL approaches reveal integration challenges: direct incorporation is either technically infeasible (e.g., in one-stage supervised CL) or degrades performance, as mixup on randomly initialized dynamic prototypes lacks meaningful impact, especially early in training. Although NC-based methods currently lead the field, this constraint highlights the need to extend SAMix to dynamic prototype-based methods like CCLIS (Li et al. 2024a).
@@ Appendix A Theoretical Proof
In this section, we prove the theory that DR loss should be used instead of contrastive loss for mixed-up samples. Our proof uses the same technique as in (Yang et al. 2022), where they demonstrate that samples from minority class should not be learned with Cross-Entroy (CE) loss. However, in this work, we extend this proof in case of mixed-up samples when learning representation.
Contrastive loss. Specifically, these samples can not be learned with current supervised contrastive loss as SupCon (Khosla et al. 2020). Indeed, the equation of SupCon loss is as follows,
@@
From equation of $\mathcal{L}_{SupCon}$, when using with mixed-up samples, since they have no positive views, $P(i)=\emptyset$ for all mixed-up samples. Therefore, $\mathcal{L}_{SupCon}$ is always equal to 0. In the context of NC with fixed prototypes, recent loss functions like FNC2 (Dang et al. 2025) (as described in Eq. (8)) can be used, but it is not efficient since it can break the optimal points. Similar to SupCon, when adapting $\mathcal{L}_{FNC^{2}}$ for mixed-up samples, $P(i)=\emptyset$ for these samples. Therefore, this loss will become as follows (for simplify, we consider this loss for a specific mixed sample $\tilde{\mathbf{x}}$ with corresponding feature $)\tilde{\mathbf{z}}=g^{t}(f^{t}(\tilde{\mathbf{x}}$))) and mixed prototype $\tilde{\mathbf{p}}$,
@@
where $A={\sum_{{z_{k}}\neq\tilde{\mathbf{z}}}{e^{\langle\tilde{\mathbf{z}}\cdot\mathbf{z}_{k}\rangle/\tau}}+\sum_{\mathbf{p}_{l}\in\mathbf{P}_{1:t-1}}{e^{\langle\tilde{\mathbf{z}}\cdot\mathbf{p}_{l}\rangle/\tau}}}$.
Here, the prototypes are fixed, only features are learnable, the gradient of $\mathcal{L}_{FNC^{2}}$ for mixed-up samples with respect to $\tilde{\mathbf{z}}$ is:
@@
$=-{\tilde{\mathbf{p}}/\tau}+{\sum_{\mathbf{z}_{k}\neq\tilde{\mathbf{z}}}{(\mathbf{z}_{k}/\tau)}\frac{e^{\langle\tilde{\mathbf{z}}\cdot\mathbf{z}_{k}\rangle}/\tau}{A}}$
$\hskip 40.00006pt+{(\mathbf{p}_{l}/\tau)}{\sum_{\mathbf{p}_{l}\in\mathbf{P}_{1:t-1}}\frac{e^{\langle\tilde{\mathbf{z}}\cdot\mathbf{p}_{l}\rangle}/\tau}{A}}$
$=-{\tilde{\mathbf{p}}/\tau}+{\sum_{\mathbf{z}_{k}\neq\tilde{\mathbf{z}}}{(\mathbf{z}_{k}/\tau)}\mathcal{F}(\mathbf{z}_{k})}$
$\hskip 40.00006pt+{\sum_{\mathbf{p}_{l}\in\mathbf{P}_{1:t-1}}{(\mathbf{p}_{l}/\tau)}{\mathcal{F}(\mathbf{p}_{l})}}$ (15)
@@
The “push” gradient $\mathbf{G}_{{push}}$ is crucial for learning classifiers. However, here, $\tilde{\mathbf{p}}$ is not learnable as it is created by mixing fixed prototypes; therefore, the “push” term is unnecessary and can even cause deviation since it does not direct to the corresponding mixed prototype as the “pull” term does. Furthermore, it can be observed that $\mathbf{G}_{{pull}}$ is dominated by $\mathbf{G}_{{push}}$, making it difficult to converge and inducing a mixed-up configuration (Fisher, Meng, and Papyan 2024). Building on these insights, we opt not to use contrastive loss such as $\mathcal{L}_{{FNC}^{2}}$ for mixed-up samples.
Dot-Regression loss. The gradient of the Dot-Regression (DR) loss is provided in (Yang et al. 2022), with a mixed-up sample $\tilde{\mathbf{x}}$ and corresponding feature $\tilde{\mathbf{z}}$, it is calculated as follows,
@@
Appendix B Neural Collapse Concept
The concept of NC is defined as follows.
where $\delta_{i,j}=1$ in case of $i=j$, and 0 otherwise.
Subsequently, the NC phenomenon can be formally characterized by the following four attributes (Papyan, Han, and Donoho 2020):
[…]
Appendix C Definitions of Distillation Losses
@@
In the loss $\mathcal{L}_{IRD}$ (Cha, Lee, and Shin 2021), $\mathbf{o}^{t}(\mathbf{x}_{i})$ is defined as:
@@
where $\kappa$ is the temperature hyperparameter, and $t\geq 1$ is the task index. For the loss $\mathcal{L}_{S-PRD}$ (Dang et al. 2025), given the prototype set $\mathbf{P}_{1:t}=\{\mathbf{p}_{s}\}_{s=1}^{S}$, which includes all prototypes used from task $1$ to task $t$. The definition of $\mathcal{L}_{S-PRD}$ is defined as,
@@
Appendix D Additional Training Details
D.1 Data Preparation
Similar to other baselines (Cha, Lee, and Shin 2021; Dang et al. 2025; Wen et al. 2024), we apply a standard set of augmentations in the CL literature for each image as follows (using Pytorch (Paszke et al. 2019) notations):
@@ D.1 Data Preparation
In addition, following (Buzzega et al. 2020; Cha, Lee, and Shin 2021; Dang et al. 2025; Wen et al. 2024), each batch is sampled equally and independently from the union of current task data and memory, eliminating the need to control old samples, avoiding sample imbalances and single-class batches that can degrade contrastive loss like FNC2 (Dang et al. 2025).
D.2 Algorithm Overview
The overview of our method is summarized in Algorithm 1.
D.3 Hyperparameter Selection
@@
Baseline. The baseline hyperparameters under evaluation include:
  • Number of initial epochs ($E_{1}$)
  • Number of epochs of $t$-th task ($E_{t\geq 2}$)
  • Learning rate ($\eta$)
  • Batch size (bsz)
  • Focusing hyperparameters ($\gamma$) for FNC2 loss
  • Temperature for the Focal Neural Collapse Contrastive (FNC2) ($\tau$)
  • Temperature for instance-wise relation distillation loss ($\mathcal{L}_{IRD}$): Following (Cha, Lee, and Shin 2021), we employ distinct temperature hyperparameters for the similarity vectors of past ($\kappa_{past}$) and current ($\kappa_{current}$) data
@@
Table 3: Search spaces of all hyperparameters. Baselines use the top set; while SAMix tunes only 4 additional hyperparameters below.
These baseline hyperparameters are tuned once for the baselines without SAMix and kept fixed when integrating SAMix.
When using SAMix. In addition to these tuned baseline values, we further tune only four hyperparameters: $M$, $\alpha$, $\upsilon$, and $\iota$, listed as:
@@
Algorithm 1 Sphere-Adaptive Mixup (SAMix) in Neural Collapse-based Continual Learning
3:for $t=1,\cdots,T$ do
4:  for batch $\mathcal{B}_{t}\sim\mathcal{D}_{t}\cup\mathcal{M}$ do
5:   $\lambda\leftarrow\text{Beta}(\alpha,\alpha)$
7:   $\mathbf{P}_{mix}\leftarrow\text{SAMix}(\mathbf{P},\lambda)$ with Eq. (6)
8:   if ploss_type is DR then
10:   else
12:   end if
13:   if t >> 1 then
15:   end if
16:   $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$
17:  end for
19:  $\theta_{prev}\leftarrow\theta$
20:end for
D.4 Training Schedules
In training. For all experiments, we choose SGD as the optimizer with momentum of 0.9 and weight decay of 0.0001. The learning rate is warmed up for 10 epochs and then decayed using the cosine schedule (Loshchilov and Hutter 2017), which restarts at the beginning of each new task.
In evaluation. A linear classifier is trained for 100 epochs using SGD with momentum 0.9 and no weight decay. The initial learning rates are 1.0, 0.1, 0.1 for Seq-Cifar-10, Seq-Cifar-100, Seq-Tiny-ImageNet, respectively, then decayed exponentially by a factor of 0.2 at epochs 60, 75, and 90.
@@ D.4 Training Schedules
Table 4: Selected baseline hyperparameters in our experiments.
Method Buffer size Dataset Hyperparameter
FC-NCCL
+ SAMix
0, 200, 500 Seq-Cifar-10 bsz: 512, $\eta$: 0.5, $\gamma$: 1, $E_{1}$: 500, $E_{t\geq 2}$: 100,
$\tau$: 0.5, $e_{0}$: 30, $\kappa_{past}$: 0.01, $\kappa_{current}$: 0.2,
$\zeta_{past}$: 0.01, $\zeta_{current}$: 0.2,
$\tau$: 0.5, $e_{0}$: 30, $\kappa_{past}$: 0.01, $\kappa_{current}$: 0.2,
$\zeta_{past}$: 0.1, $\zeta_{current}$: 0.2,
$\tau$: 0.5, $e_{0}$: 20, $\kappa_{past}$: 0.1, $\kappa_{current}$: 0.1,
$\zeta_{past}$: 0.1, $\zeta_{current}$: 0.2,
TA-NCCL
+ SAMix
0, 200, 500 Seq-Cifar-10 bsz: 512, $\eta$: 0.5, $E_{1}$: 500, $E_{t\geq 2}$: 100,
$\tau$: 0.5, $e_{0}$: 30, $\kappa_{past}$: 0.01, $\kappa_{current}$: 0.2,
$\zeta_{past}$: 0.01, $\zeta_{current}$: 0.2,
$\tau$: 0.5, $e_{0}$: 30, $\kappa_{past}$: 0.01, $\kappa_{current}$: 0.2,
$\zeta_{past}$: 0.1, $\zeta_{current}$: 0.2,
$\tau$: 0.5, $e_{0}$: 20, $\kappa_{past}$: 0.1, $\kappa_{current}$: 0.1,
$\zeta_{past}$: 0.1, $\zeta_{current}$: 0.2,
Table 5: Selected hyperparameters when integrating SAMix. Differences from the baselines are shown in bold, only 4 additional hyperparameters are tuned.
Appendix E Additional Results
E.1 Average forgetting results
The Average Forgetting metric (Chaudhry et al. 2018), used in this work to quantify model forgetting, is defined as:
@@ E.1 Average forgetting results
Buffer Dataset Venue Seq-Cifar-100 Seq-Tiny-ImageNet Seq-Cifar-10
@@
Table 6: Average forgetting (lower is better) across 5 independent trials from 5 different seeds: Comparing our method against all baselines in continual learning. (Ours: TA-NCCL method)
E.2 Calibration Error and Overconfidence
@@
0 FC-NCCL (Dang et al. 2025) 0.160 0.107 74.25 0.130 0.050 69.70 0.206 0.167 88.30
FC-NCCL (Dang et al. 2025)(+ SAMix) 0.141 $\blacktriangle$ 0.082 $\blacktriangle$ 77.00 $\blacktriangle$ 0.088 $\blacktriangle$ 0.021 $\blacktriangle$ 74.40 $\blacktriangle$ 0.177 $\blacktriangle$ 0.133 $\blacktriangle$ 85.60 $\triangledown$
@@
Ours(+ SAMix) 0.106 $\blacktriangle$ 0.005 $\blacktriangle$ 78.55 $\blacktriangle$ 0.146 $\blacktriangle$ 0.069 $\blacktriangle$ 63.30 $\blacktriangle$ 0.152 $\triangledown$ 0.099 $\triangledown$ 85.15 $\blacktriangle$
200 FC-NCCL (Dang et al. 2025) 0.162 0.109 73.85 0.170 0.073 69.70 0.305 0.252 95.70
FC-NCCL (Dang et al. 2025)(+ SAMix) 0.139 $\blacktriangle$ 0.046 $\blacktriangle$ 77.30 $\blacktriangle$ 0.095 $\blacktriangle$ 0.033 $\blacktriangle$ 73.40 $\blacktriangle$ 0.312 $\triangledown$ 0.273 $\triangledown$ 89.30 $\triangledown$
@@
Ours(+ SAMix) 0.098 $\blacktriangle$ 0.017 $\blacktriangle$ 80.15 $\blacktriangle$ 0.146 $\blacktriangle$ 0.069 $\blacktriangle$ 64.70 $\blacktriangle$ 0.219 $\triangledown$ 0.172 $\triangledown$ 91.25 $\blacktriangle$
500 FC-NCCL (Dang et al. 2025) 0.231 0.164 71.95 0.098 0.040 63.20 0.334 0.290 96.40
FC-NCCL (Dang et al. 2025)(+ SAMix) 0.132 $\blacktriangle$ 0.092 $\blacktriangle$ 76.50 $\blacktriangle$ 0.079 $\blacktriangle$ 0.031 $\blacktriangle$ 64.20 $\blacktriangle$ 0.267 $\blacktriangle$ 0.226 $\blacktriangle$ 93.60 $\triangledown$
@@
Ours(+ SAMix) 0.078 $\blacktriangle$ 0.015 $\blacktriangle$ 80.05 $\blacktriangle$ 0.115 $\blacktriangle$ 0.052 $\blacktriangle$ 56.80 $\blacktriangle$ 0.150 $\triangledown$ 0.122 $\triangledown$ 92.55 $\blacktriangle$
Table 7: AECE, AOE, and last task accuracy $(A_{T,T})$ across different NC-based methods. $\blacktriangle$$\triangledown$ indicate methods applying SAMix outperforms/underperforms compared to methods without SAMix. All results indicate that SAMix is highly effective on large, complex datasets such as Seq-Cifar-100 and Seq-Tiny-ImageNet but largely ineffective on smaller datasets like Seq-Cifar-10.
@@
E.3 Additional Ablation Studies
Extra time when using SAMix.
Although our method increase batch size (see Sec. 3.4), mixed samples are used only in the plasticity loss (DR), not in distillation or contrastive losses like FNC2 to avoid disrupting the mixed-up configuration (see Appendix proof). Thus, mixed samples are not used as regular inputs, keeping the impact on training time minimal. Table 8 shows the extra time when applying SAMix to FC-NCCL and TA-NCCL on Seq-Cifar-100 (CIL setting). Across buffer sizes, the additional time is a good trade-off, justified by significant performance gains. As in the discussion of (Chen et al. 2025), other SOTA CL methods such as L2P (Wang et al. 2022b) and DualPrompt (Wang et al. 2022a) increase training time by $37\%$ and $32\%$, respectively, for performance gains of only $2.5\%$ and $1.4\%$.
Performance with reduced batch size.
To evaluate SAMix under smaller batches, we run TA-NCCL on Seq-Cifar-10 (Class-IL setting) with a batch size of 256 with different buffer sizes. As shown in Table 9, even without memory, SAMix with half the batch size outperforms the SOTA method CCLIS (Li et al. 2024a), which uses a batch size of 512 and memory size of 200. This highlights the strong adaptability of our method under constrained settings.
Buffer 0 200 500
Overhead Gain Overhead Gain Overhead Gain
FC-NCCL 14.1 7.6 16.5 7.9 19.8 7.6
TA-NCCL 22.3 10.9 21.7 10.2 28.1 14.1
Table 8: Extra time (overhead) and performance gain ($\%$) with SAMix on Seq-Cifar-100 (CIL setting).
Buffer Method Batch size AA
0 TA-NCCL 512 44.41
TA-NCCL 256 42.77
200 CCLIS 512 42.39
TA-NCCL 512 46.95
TA-NCCL 256 45.21
Table 9: SAMix results with varying batch sizes on Seq-Cifar-100 (CIL setting). Bold indicates comparison purposes.

### v2

@@ 5 Conclusion
However, a limitation exists: while SAMix improves performance and calibration, it is tailored to fixed prototype-based CL. Our experiments with dynamic prototype methods and traditional one-stage CL approaches reveal integration challenges: direct incorporation is either technically infeasible (e.g., in one-stage supervised CL) or degrades performance, as mixup on randomly initialized dynamic prototypes lacks meaningful impact, especially early in training. Although NC-based methods currently lead the field, this constraint highlights the need to extend SAMix to dynamic prototype-based methods like CCLIS [li_cl_important_sampling_2024]. Moreover, like other NC-based CL methods, SAMix assumes prior knowledge of the total number of classes to predefine prototypes. Recent work [markou2024guidingneuralcollapseoptimising] on dynamically aligning classifiers to the nearest simplex ETF suggests that such structures need not be fixed in advance. Extending SAMix to incrementally grow the simplex as new classes emerge in CL is a promising direction for future work.
[…]
Acknowledgements This work is partially supported by the ANR-21-ASRO-0003 ROV-Chasseur project.
@@ Appendix A Theoretical Proof
In this section, we prove the theory that DR loss should be used instead of contrastive loss for mixed-up samples. Our proof uses the same technique as in [yang_induce_nc_2022], where they demonstrate that samples from minority class should not be learned with Cross-Entroy (CE) loss. However, in this work, we extend this proof in case of mixed-up samples when learning representation. Contrastive loss. Specifically, these samples can not be learned with current supervised contrastive loss as SupCon [khosla2021supcon]. Indeed, the equation of SupCon loss is as follows,
@@
From equation of $\mathcal{L}_{SupCon}$, when using with mixed-up samples, since they have no positive views, $P(i)=\emptyset$ for all mixed-up samples. Therefore, $\mathcal{L}_{SupCon}$ is always equal to 0. In the context of NC with fixed prototypes, recent loss functions like FNC2 [anh_wacv_2025] (as described in Eq. (10)) can be used, but it is not efficient since it can break the optimal points. Similar to SupCon, when adapting $\mathcal{L}_{FNC^{2}}$ for mixed-up samples, $P(i)=\emptyset$ for these samples. Therefore, this loss will become as follows (for simplify, we consider this loss for a specific mixed sample $\tilde{\mathbf{x}}$ with corresponding feature $)\tilde{\mathbf{z}}=g^{t}(f^{t}(\tilde{\mathbf{x}}$))) and mixed prototype $\tilde{\mathbf{p}}$,
@@
where $A={\sum_{{z_{k}}\neq\tilde{\mathbf{z}}}{e^{\langle\tilde{\mathbf{z}}\cdot\mathbf{z}_{k}\rangle/\tau}}+\sum_{\mathbf{p}_{l}\in\mathbf{P}_{1:t-1}}{e^{\langle\tilde{\mathbf{z}}\cdot\mathbf{p}_{l}\rangle/\tau}}}$. Here, the prototypes are fixed, only features are learnable, the gradient of $\mathcal{L}_{FNC^{2}}$ for mixed-up samples with respect to $\tilde{\mathbf{z}}$ is:
@@
$=-{\tilde{\mathbf{p}}/\tau}+{\sum_{\mathbf{z}_{k}\neq\tilde{\mathbf{z}}}{(\mathbf{z}_{k}/\tau)}\frac{e^{\langle\tilde{\mathbf{z}}\cdot\mathbf{z}_{k}\rangle}/\tau}{A}}+{(\mathbf{p}_{l}/\tau)}{\sum_{\mathbf{p}_{l}\in\mathbf{P}_{1:t-1}}\frac{e^{\langle\tilde{\mathbf{z}}\cdot\mathbf{p}_{l}\rangle}/\tau}{A}}$
$=-{\tilde{\mathbf{p}}/\tau}+{\sum_{\mathbf{z}_{k}\neq\tilde{\mathbf{z}}}{(\mathbf{z}_{k}/\tau)}\mathcal{F}(\mathbf{z}_{k})}+{\sum_{\mathbf{p}_{l}\in\mathbf{P}_{1:t-1}}{(\mathbf{p}_{l}/\tau)}{\mathcal{F}(\mathbf{p}_{l})}}$ (17)
@@
Appendix B Definitions of Distillation Losses
@@
In the loss $\mathcal{L}_{IRD}$ [co2l], $\mathbf{o}^{t}(\mathbf{x}_{i})$ is defined as:
@@
where $\kappa$ is the temperature hyperparameter, and $t\geq 1$ is the task index. For the loss $\mathcal{L}_{S-PRD}$ [anh_wacv_2025], given the prototype set $\mathbf{P}_{1:t}=\{\mathbf{p}_{s}\}_{s=1}^{S}$, which includes all prototypes used from task $1$ to task $t$. The definition of $\mathcal{L}_{S-PRD}$ is defined as,
@@
Appendix C Additional Training Details
C.1 Data Preparation
Similar to other baselines [co2l, anh_wacv_2025, wen2024provablecontrastivecontinuallearning], we apply a standard set of augmentations in the CL literature for each image as follows (using Pytorch [paszke2019pytorchimperativestylehighperformance] notations):
@@ C.1 Data Preparation
In addition, following [buzzega2020darkexperiencegeneralcontinual, co2l, anh_wacv_2025, wen2024provablecontrastivecontinuallearning], each batch is sampled equally and independently from the union of current task data and memory, eliminating the need to control old samples, avoiding sample imbalances and single-class batches that can degrade contrastive loss like FNC2 [anh_wacv_2025].
C.2 Hyperparameter Selection
@@
Table 5: Search spaces of all hyperparameters. Baselines use the top set; while SAMix tunes only 4 additional hyperparameters below
@@
Baseline. The baseline hyperparameters under evaluation include:
  • Number of initial epochs ($E_{1}$)
  • Number of epochs of $t$-th task ($E_{t\geq 2}$)
  • Learning rate ($\eta$)
  • Batch size (bsz)
  • Focusing hyperparameters ($\gamma$) for FNC2 loss
  • Temperature for the Focal Neural Collapse Contrastive (FNC2) ($\tau$)
  • Temperature for instance-wise relation distillation loss ($\mathcal{L}_{IRD}$): Following [co2l], we employ distinct temperature hyperparameters for the similarity vectors of past ($\kappa_{past}$) and current ($\kappa_{current}$) data
These baseline hyperparameters are tuned once for the baselines without SAMix and kept fixed when integrating SAMix. When using SAMix. In addition to these tuned baseline values, we further tune only four hyperparameters: $M$, $\alpha$, $\upsilon$, and $\iota$, listed as:
@@
C.3 Training Schedules
In training. For all experiments, we choose SGD as the optimizer with momentum of 0.9 and weight decay of 0.0001. The learning rate is warmed up for 10 epochs and then decayed using the cosine schedule [loshchilov2017sgdrstochasticgradientdescent], which restarts at the beginning of each new task. In evaluation. A linear classifier is trained for 100 epochs using SGD with momentum 0.9 and no weight decay. The initial learning rates are 1.0, 0.1, 0.1 for Seq-Cifar-10, Seq-Cifar-100, Seq-Tiny-ImageNet, respectively, then decayed exponentially by a factor of 0.2 at epochs 60, 75, and 90.
Table 6: Selected hyperparameters when integrating SAMix. Differences from the baselines are shown in bold, only 4 additional hyperparameters are tuned
Method Buffer size Dataset Hyperparameter
FC-NCCL + SAMix 0, 200, 500 Seq-Cifar-10 bsz: 512, $\eta$: 0.5, $\gamma$: 1, $E_{1}$: 500, $E_{t\geq 2}$: 100,
$\tau$: 0.5, $e_{0}$: 30, $\kappa_{past}$: 0.01, $\kappa_{current}$: 0.2,
$\zeta_{past}$: 0.01, $\zeta_{current}$: 0.2,
$\tau$: 0.5, $e_{0}$: 30, $\kappa_{past}$: 0.01, $\kappa_{current}$: 0.2,
$\zeta_{past}$: 0.1, $\zeta_{current}$: 0.2,
$\tau$: 0.5, $e_{0}$: 20, $\kappa_{past}$: 0.1, $\kappa_{current}$: 0.1,
$\zeta_{past}$: 0.1, $\zeta_{current}$: 0.2,
TA-NCCL + SAMix 0, 200, 500 Seq-Cifar-10 bsz: 512, $\eta$: 0.5, $E_{1}$: 500, $E_{t\geq 2}$: 100,
$\tau$: 0.5, $e_{0}$: 30, $\kappa_{past}$: 0.01, $\kappa_{current}$: 0.2,
$\zeta_{past}$: 0.01, $\zeta_{current}$: 0.2,
$\tau$: 0.5, $e_{0}$: 30, $\kappa_{past}$: 0.01, $\kappa_{current}$: 0.2,
$\zeta_{past}$: 0.1, $\zeta_{current}$: 0.2,
$\tau$: 0.5, $e_{0}$: 20, $\kappa_{past}$: 0.1, $\kappa_{current}$: 0.1,
$\zeta_{past}$: 0.1, $\zeta_{current}$: 0.2,
Table 7: Selected baseline hyperparameters in our experiments
@@ C.3 Training Schedules
Appendix D Additional Results
D.1 Average forgetting results
The Average Forgetting metric [chaudhry_2018], used in this work to quantify model forgetting, is defined as:
@@ D.1 Average forgetting results
Table 8: Average forgetting (lower is better) across 5 independent trials from 5 different seeds: Comparing our method against all baselines in continual learning (Ours: TA-NCCL method)
Buffer Dataset Seq-Cifar-100 Seq-Tiny-ImageNet Seq-Cifar-10
@@
D.2 Calibration Error and Overconfidence
Table 9: AECE, AOE, and last task accuracy $(A_{T,T})$ across different NC-based methods. $\blacktriangle$$\triangledown$ indicate methods applying SAMix outperforms/underperforms compared to methods without SAMix. All results indicate that SAMix is highly effective on large, complex datasets such as Seq-Cifar-100 and Seq-Tiny-ImageNet but largely ineffective on smaller datasets like Seq-Cifar-10
@@ D.2 Calibration Error and Overconfidence
0 FC-NCCL [anh_wacv_2025] 0.160 0.107 74.25 0.130 0.050 69.70 0.206 0.167 88.30
FC-NCCL + SAMix 0.141 $\blacktriangle$ 0.082 $\blacktriangle$ 77.00 $\blacktriangle$ 0.088 $\blacktriangle$ 0.021 $\blacktriangle$ 74.40 $\blacktriangle$ 0.177 $\blacktriangle$ 0.133 $\blacktriangle$ 85.60 $\triangledown$
@@
Ours + SAMix 0.106 $\blacktriangle$ 0.005 $\blacktriangle$ 78.55 $\blacktriangle$ 0.146 $\blacktriangle$ 0.069 $\blacktriangle$ 63.30 $\blacktriangle$ 0.152 $\triangledown$ 0.099 $\triangledown$ 85.15 $\blacktriangle$
200 FC-NCCL [anh_wacv_2025] 0.162 0.109 73.85 0.170 0.073 69.70 0.305 0.252 95.70
FC-NCCL + SAMix 0.139 $\blacktriangle$ 0.046 $\blacktriangle$ 77.30 $\blacktriangle$ 0.095 $\blacktriangle$ 0.033 $\blacktriangle$ 73.40 $\blacktriangle$ 0.312 $\triangledown$ 0.273 $\triangledown$ 89.30 $\triangledown$
@@
Ours + SAMix 0.098 $\blacktriangle$ 0.017 $\blacktriangle$ 80.15 $\blacktriangle$ 0.146 $\blacktriangle$ 0.069 $\blacktriangle$ 64.70 $\blacktriangle$ 0.219 $\triangledown$ 0.172 $\triangledown$ 91.25 $\blacktriangle$
500 FC-NCCL [anh_wacv_2025] 0.231 0.164 71.95 0.098 0.040 63.20 0.334 0.290 96.40
FC-NCCL + SAMix 0.132 $\blacktriangle$ 0.092 $\blacktriangle$ 76.50 $\blacktriangle$ 0.079 $\blacktriangle$ 0.031 $\blacktriangle$ 64.20 $\blacktriangle$ 0.267 $\blacktriangle$ 0.226 $\blacktriangle$ 93.60 $\triangledown$
@@
Ours + SAMix 0.078 $\blacktriangle$ 0.015 $\blacktriangle$ 80.05 $\blacktriangle$ 0.115 $\blacktriangle$ 0.052 $\blacktriangle$ 56.80 $\blacktriangle$ 0.150 $\triangledown$ 0.122 $\triangledown$ 92.55 $\blacktriangle$